OpenStudy (anonymous):

The radius of a cylinder of a circular oil slick expands at a rate of 2m/min. How fast is the area of the oil slick increasing when the radius is 25m?

OpenStudy (anonymous):

\[ A(t)=\pi [r(t)]^2 \]They're both functions of \(t\). You an differentiate with chain rule: \[ A'(t)=\pi 2r(t)r'(t) \]

OpenStudy (anonymous):

You can plug in the values they gave you. \(r=25\) and \(r'=2\): \[ A'(t) = 2\pi(25)(2) \]